OpenStudy (anonymous):

how many atoms of gold are in a 5 gram sample of pure gold?

OpenStudy (koikkara):

Hai Sweet Friend, @eng0ling WelcoMe To OpeN studY !! There are 197 grams per mole for gold, so 5 grams / 197 = .0254 moles There are 6.022 x 10^23 molecules per mole, so multiply by Avigadro's number, .0254 x 6.022 x 10^23 = 1.5284 x 10^22 Hope, if U Are Satisfied with this Answer, Please Close This Question ! Thank U !! Keep In Touch, with Open study !!

OpenStudy (abb0t):

5 grams of Au \(\sf \times \frac{mol}{196.96~g~Au}\) = moles of Au. Now use avogadro's number: \(\sf \color{red}{6.022 \times 10^{23}\frac{atom}{mol}}\)